# MAVERIX: Multimodal Audio-Visual Evaluation Reasoning IndeX

### v1

@@ 5 Conclusion
We are releasing MAVERIX v1.0, our proposed benchmark dataset for multimodal video-language understanding. The benchmark dataset is provided as a single JSON file for ease of use and straightforward integration with existing benchmarking pipelines. For each video, the dataset includes metadata and contains multiple-choice and open-ended questions covering multiple tasks from our proposed task suite. Each task is accompanied by a set of questions designed to evaluate visual reasoning, situational awareness, and cross-modal understanding. Relevant timestamps are provided to allow precise video trimming.
@@
We provide all 700 video clips used in our benchmark, and users can download the corresponding videos from the Ego4D website after reviewing and accepting the Ego4D license agreement. We also provide sample videos with annotations from MAVERIX. All materials are available through the project page upon acceptance.
@@
    • vid/: Contains all videos for evaluations without audio modality.
@@
    • video_utils.py: A script for dataloading, processing, and evaluation functionalities.
@@
    • MAVERIXXX_dev: Individual setting files specifying the test settings for different modes. The currently supported modes are: audio_only, video_only, text_only, video_text, and video_text_audio.
    • Remark: Except for the open-sourced models listed in the paper or available through LMMs-Eval[58, 26], all other experiments require access to proprietary models.
@@
Appendix B Rationale and limitations
@@
Appendix C Qualitative Comparisons by Modalities
We conduct a comprehensive analysis of the performance of Gemini 1.5 Pro, providing a breakdown across each custom-defined split and input modality to better understand its strengths and limitations throughout the benchmark. Specifically, we evaluate the model under different input configurations like Audio Only, Video Only, and Audio+++Video, as well as across diverse content-based splits covering topics, agentic task types, audio characteristics, and difficulty levels. The corresponding results are shown in Fig. C.1 for modality comparison, Fig. C.2 for topic-wise performance, Fig. C.3 for agentic task types, Fig. C.4 for audio type variation, and Fig. C.5 for difficulty levels. This detailed analysis highlights how well Gemini 1.5 Pro generalizes across different challenge dimensions and input conditions.
Figure C.1: Evaluation of multimodal abilities by Categories.
Figure C.2: Evaluation of multimodal abilities by Topics.
Figure C.3: Evaluation of multimodal abilities by Agentic Tasks.
Figure C.4: Evaluation of multimodal abilities by Audio Types.
Figure C.5: Evaluation of multimodal abilities by the level of difficulties.
Appendix D Additional experiments
@@
D.1 Standardized Prompt Design
To ensure fairness in evaluation, we adopt a standardized prompt format across all models. This eliminates potential biases introduced by prompt engineering and ensures that differences in performance are attributed to model capabilities rather than variations in input phrasing. For tasks requiring multiple modalities, the prompts incorporate the modalities in the inputs. In multiple-choice settings, the prompt is presented as follows: {mdframed} [images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available. [Subtitles] Select the best answer to the following multiple-choice question based on the video and the subtitles. Respond with only the letter (A, B, C, D, E, F, G, or H) of the correct option. Respond with only the letter (A, B, C, or D) of the correct option. [Question] The best answer is:
@@ D.1 Standardized Prompt Design
D.2 Model Performance by Difficulties
We evaluate model performance across different difficulty levels to assess robustness and generalization from simpler to more complex scenarios. The results are presented in Figure D.1, which illustrates how the model’s accuracy declines as task complexity increases and showcases its relative performance compared to human participants.
Figure D.1: Probed model performance across different difficulty levels. This analysis illustrates how the models’ accuracy changes with increasing task complexity and their relative positions against humans.
D.3 Human Performance by Categories
We report performance across distinct agentic task categories to showcase the strengths and weaknesses of the human study participants within the defined contexts. We note that these statistics may be biased due to the composition of the participant pool, and further details on the recruitment process can be found in Sec. E.
@@ D.3 Human Performance by Categories
Visual Social 46.15 66.67 54.87
Sports 20.00 100.00 51.43
Information Querying 53.85 71.43 62.15
Sentiment 20.00 73.33 47.71
Shopping 35.71 75.00 57.68
Gaming 37.50 100.00 58.24
Overall 35.20 81.07 55.68
Audio Social 16.67 40.00 24.17
Sports 20.00 42.86 33.57
Information Querying 23.08 36.67 33.27
Sentiment 21.43 40.00 35.36
Shopping 23.08 60.00 44.10
Gaming 25.00 25.00 25.00
Overall 21.88 40.42 32.58
Both Social 66.67 100.00 79.37
Sports 42.86 100.00 80.95
Information Querying 73.33 89.66 81.00
Sentiment 86.67 93.33 89.17
Shopping 60.00 100.00 80.00
Gaming 57.14 87.50 71.13
Overall 64.78 95.08 80.94
Table D.1: Worst, Best, and Average human performance (%) across tasks and modalities.
D.4 Quantitative GPT Evaluations
To further assess model performance beyond multiple-choice accuracy, we conduct a quantitative evaluation of the open-ended question responses using GPT-based scoring. Specifically, as discussed in Sec. 2.2, we prompt GPT-4o to act as an automated grader, evaluating the correctness and relevance of the model’s free-form answers based on predefined guidelines consistent with the task objectives.
@@ D.4 Quantitative GPT Evaluations
We report the GPT-assigned scores across the full benchmark as well as per agentic task category, enabling a more granular understanding of the model’s strengths and weaknesses in open-ended scenarios. These findings complement the multiple-choice evaluations and expose areas where the model demonstrates strong language generation and areas where further improvement is needed. However, we note that these scores may be biased due to limitations in GPT-4o’s own reasoning and evaluation capabilities [28], which can introduce alignment artifacts or systematic preferences.
Models Overall
Human (Video) 1.03
Human (Audio) 0.65
Human (Video+Audio) 2.50
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.896 1.055
Qwen2.5-VL 1.651 1.929
LLaVA-OneVision 1.774 1.124
DeepSeek-VL2-Tiny 0.839 0.870
DeepSeek-VL2-Small 1.128 1.471
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.327 2.308
o1 2.697 3.075
Claude 3.5 Sonnet 1.964 2.568
Grok2-Vision 2.193 2.765
Nova-Lite 1.444 1.859
Nova-Pro 1.998 2.364
Gemini 1.5 Pro 2.170 2.538
Open source MLLMs w/ Full-Length Videos
InternVL2 0.895 1.102
Qwen2.5-VL 1.489 1.830
LLaVA-OneVision 1.471 1.666
DeepSeek-VL2-Tiny 0.675 0.808
DeepSeek-VL2-Small 0.839 1.295
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.061 2.148
o1 2.484 2.983
Claude 3.5 Sonnet 0.017 0.401
Grok2-Vision 1.935 2.622
Nova-Lite 1.653 1.708
Nova-Pro 1.615 2.012
Gemini 1.5 Pro 1.956 2.318
Table D.2: GPT-4o evaluation of consistency score.
Models Overall
Human (Video) 1.02
Human (Audio) 0.58
Human (Video+Audio) 2.71
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.129 1.007
Qwen2.5-VL 1.404 1.702
LLaVA-OneVision 1.168 1.005
DeepSeek-VL2-Tiny 0.562 0.614
DeepSeek-VL2-Small 1.087 1.851
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.277 2.103
o1 2.484 2.803
Claude 3.5 Sonnet 2.122 2.283
Grok2-Vision 1.901 2.434
Nova-Lite 1.287 1.602
Nova-Pro 1.518 1.741
Gemini 1.5 Pro 1.970 2.299
Open source MLLMs w/ Full-Length Videos
InternVL2 1.333 0.975
Qwen2.5-VL 1.275 1.660
LLaVA-OneVision 1.277 1.572
DeepSeek-VL2-Tiny 0.687 1.168
DeepSeek-VL2-Small 0.873 1.268
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.277 2.103
o1 2.484 2.803
Claude 3.5 Sonnet 0.319 0.415
Grok2-Vision 1.693 2.189
Nova-Lite 1.196 1.481
Nova-Pro 1.384 1.678
Gemini 1.5 Pro 1.808 2.136
Table D.3: GPT-4o evaluation of correctness score.
Models Overall
Human (Video) 0.90
Human (Audio) 0.52
Human (Video+Audio) 2.55
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.586 1.120
Qwen2.5-VL 1.746 2.033
LLaVA-OneVision 0.953 1.090
DeepSeek-VL2-Tiny 0.653 0.68
DeepSeek-VL2-Small 1.162 1.493
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.284 2.220
o1 3.012 3.236
Claude 3.5 Sonnet 1.984 2.613
Grok2-Vision 2.454 2.968
Nova-Lite 1.339 1.701
Nova-Pro 1.812 1.841
Gemini 1.5 Pro 2.374 2.716
Open source MLLMs w/ Full-Length Videos
InternVL2 1.107 1.159
Qwen2.5-VL 1.559 1.996
LLaVA-OneVision 1.40 1.679
DeepSeek-VL2-Tiny 0.839 0.953
DeepSeek-VL2-Small 0.970 1.382
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.032 1.967
o1 2.880 3.159
Claude 3.5 Sonnet 0.146 0.494
Grok2-Vision 2.237 2.792
Nova-Lite 1.273 1.562
Nova-Pro 1.567 1.832
Gemini 1.5 Pro 2.288 2.618
Table D.4: GPT-4o evaluation of detailed orientation score.
Models Overall
Human (Video) 1.26
Human (Audio) 0.58
Human (Video+Audio) 2.71
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.129 1.393
Qwen2.5-VL 2.007 2.249
LLaVA-OneVision 1.168 1.311
DeepSeek-VL2-Tiny 0.839 0.870
DeepSeek-VL2-Small 1.505 1.851
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.586 2.499
o1 3.032 3.291
Claude 3.5 Sonnet 2.122 2.734
Grok2-Vision 2.468 2.960
Nova-Lite 1.742 2.039
Nova-Pro 2.046 2.238
Gemini 1.5 Pro 2.435 2.801
Open source MLLMs w/ Full-Length Videos
InternVL2 1.332 1.366
Qwen2.5-VL 1.818 2.187
LLaVA-OneVision 1.725 1.995
DeepSeek-VL2-Tiny 1.067 1.168
DeepSeek-VL2-Small 1.292 1.655
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.274 2.250
o1 2.912 3.214
Claude 3.5 Sonnet 0.319 0.533
Grok2-Vision 2.240 2.751
Nova-Lite 1.653 1.936
Nova-Pro 1.915 2.2348
Gemini 1.5 Pro 2.257 2.618
Table D.5: GPT-4o evaluation of context score.
D.5 Quantitative Evaluations by Categories
To provide a comprehensive understanding of model performance across the diverse challenge dimensions of the benchmark, we report quantitative results for all probed models across each defined split. This includes evaluations broken down by agentic task categories, difficulty levels, audio characteristics, topics, and input modalities.
@@ D.5 Quantitative Evaluations by Categories
GPT-4o o1 Gemini 1.5 Pro Nova Pro Grok2 Claude 3.5 Sonnet
Difficulty Easy 69.8 79.3 69.5 51.5 65.1 68.9
Medium 67.0 76.0 69.4 47.5 61.2 64.6
Hard 49.6 67.4 49.6 34.1 47.4 44.4
Agentic Tasks Social 64.6 69.7 62.6 49.5 60.6 66.7
Sentiment 56.8 74.1 66.7 46.9 51.9 51.9
Egocentric Agent 62.7 69.5 57.6 40.7 54.2 64.4
Information Querying 67.5 75.5 69.4 47.2 62.6 64.9
Sports 66.4 79.0 64.7 40.3 58.8 56.3
Gaming 64.1 72.8 58.3 54.4 52.4 59.2
Shopping 68.3 85.7 74.6 47.6 73.0 73.0
Audio Type Natural Sound 57.4 66.2 54.4 35.3 50.0 48.5
Speech 67.7 76.7 69.3 44.8 61.7 65.0
Music 57.7 61.5 69.2 38.5 57.7 57.7
Artificial Sound 70.0 78.0 56.0 58.0 54.0 60.0
Mixed Sounds 63.4 78.2 66.0 52.3 63.0 64.9
Topics Humanities and Society 71.4 78.9 77.4 51.1 66.9 72.2
Geography and Travel 56.2 68.8 50.0 37.5 43.8 43.8
Technology and Gaming 65.6 71.8 61.1 50.4 53.4 57.3
Science and Knowledge 60.7 64.3 53.6 28.6 50.0 53.6
Movies, TV and Animations 55.2 68.7 55.2 40.3 43.3 55.2
Arts and Performance 51.7 65.5 58.6 41.4 58.6 75.9
Pets and Animals 62.2 68.9 64.4 48.9 62.2 57.8
Business and Commerce 72.1 86.0 79.1 60.5 76.7 75.6
Life and Practical Skills 74.0 79.8 77.9 51.9 71.2 69.2
Sports and Adventure 63.5 80.2 65.1 46.0 57.1 57.9
Social Trends and Reactions 55.6 77.8 57.4 33.3 61.1 59.3
Vehicles and Transportation 88.9 100.0 55.6 55.6 55.6 55.6
Low Quality and Extended Content 54.2 58.3 37.5 16.7 50.0 54.2
Multimodal Abilities Visual Reasoning 59.2 73.2 64.5 42.1 57.5 59.6
Emotional Inference 60.5 69.8 62.8 46.5 53.5 48.8
Situational Reasoning 65.0 80.0 70.0 40.0 75.0 75.0
Causal Reasoning 65.7 73.1 67.2 50.7 58.2 68.7
Factual Recall 68.9 79.2 68.2 49.2 63.2 65.0
Spatial Reasoning 62.1 62.1 58.6 41.4 44.8 55.2
Relationship 75.0 81.2 62.5 50.0 81.2 75.0
Temporal Distance 66.7 66.7 58.3 58.3 50.0 66.7
Overall 65.4 75.9 66.3 46.9 60.6 63.1
@@
Table D.6: Performance comparison across different models and categories on truncated videos.
GPT-4o o1 Gemini 1.5 Pro Nova Pro Grok2 Claude 3.5 Sonnet
Difficulty Easy 64.8 73.0 69.4 58.2 59.2 17.4
Medium 62.2 72.4 68.0 58.1 55.2 14.8
Hard 46.5 65.4 55.1 41.7 53.5 17.3
Agentic Tasks Social 66.7 65.6 59.4 63.5 62.5 9.4
Sentiment 54.3 67.9 64.2 46.9 55.6 24.7
Egocentric Agent 55.3 61.7 55.3 19.1 34.0 14.9
Information Querying 63.9 75.4 68.3 60.7 57.9 17.5
Sports 55.3 69.9 66.0 55.3 59.2 14.6
Gaming 60.0 71.8 64.7 48.2 48.2 22.4
Shopping 60.4 75.7 76.6 64.0 62.2 10.8
Audio Type Natural Sound 70.5 79.8 61.5 53.8 43.1 30.8
Speech 56.2 81.2 70.3 57.3 60.7 14.4
Music 55.3 70.2 56.0 48.0 56.0 16.0
Artificial Sound 51.9 59.3 58.0 48.0 48.0 28.0
Mixed Sounds 58.2 59.7 63.8 55.0 54.6 12.8
Topics Humanities and Society 57.1 67.9 79.1 66.7 64.3 27.1
Geography and Travel 60.0 68.9 68.8 56.2 56.2 18.8
Technology and Gaming 61.4 72.9 62.3 50.9 50.0 17.5
Science and Knowledge 69.4 75.3 63.0 55.6 51.9 11.1
Movies, TV and Animations 55.0 68.8 59.7 52.2 46.3 11.9
Arts and Performance 57.7 75.0 57.1 46.4 57.1 17.9
Pets and Animals 44.4 88.9 60.0 57.8 57.8 17.8
Business and Commerce 62.5 62.5 77.1 65.7 67.1 12.9
Life and Practical Skills 57.8 68.6 64.7 57.6 62.4 8.2
Sports and Adventure 43.9 65.9 69.7 52.3 55.0 16.5
Social Trends and Reactions 72.2 77.8 57.7 51.9 51.9 15.4
Vehicles and Transportation 53.4 63.8 44.4 44.4 33.3 0.0
Low Quality and Extended Content 63.3 73.9 50.0 20.8 50.0 8.3
Multimodal Abilities Visual Reasoning 62.5 66.7 64.2 51.5 54.4 14.7
Emotional Inference 68.8 75.0 58.5 53.7 43.9 19.5
Situational Reasoning 81.8 90.9 66.7 55.6 77.8 16.7
Causal Reasoning 53.8 56.9 65.5 55.2 55.2 25.9
Factual Recall 63.8 71.9 69.0 58.1 58.8 15.9
Spatial Reasoning 52.0 60.0 45.8 41.7 41.7 16.7
Relationship 66.0 82.0 75.0 68.8 68.8 0.0
Temporal Distance 56.4 73.9 81.8 54.5 45.5 18.2
Overall 60.6 71.5 66.5 55.5 56.5 16.3
Table D.7: Performance comparison across different models and categories on full-length videos.
D.6 Common Error Modes

### v2

@@ Conclusion
We are releasing MAVERIX v1.0, our proposed benchmark dataset for multimodal video-language understanding, built with support to a common benchmark platform LMMs-Eval (Li* et al. 2024). The benchmark dataset is provided as a single JSON file for ease of use and straightforward integration with existing benchmarking pipelines. For each video, the dataset includes metadata and contains multiple-choice and open-ended questions covering multiple tasks from our proposed task suite. Each task is accompanied by a set of questions designed to evaluate visual reasoning, situational awareness, and cross-modal understanding. Relevant timestamps are provided to allow precise video trimming.
@@
We will provide all 1400 video clips used in our benchmark through HuggingFace (hug ), with both localized and full length videos, and users can download the corresponding videos from the Ego4D website after reviewing and accepting the Ego4D license agreement. We also provide sample videos with annotations from MAVERIX. All materials are available through the project page upon acceptance. The code snippets providing main input prompts, response aggregation, judge response generations are provided as part of the supplement materials, with unit test functions for testing data corruption included.
@@
    • vid/: Contains all videos for evaluations with both video and audio modality.
    • audio/: Contains all audio files for evaluations with audio modality.
@@
    • maverix_benchmark.py: A script for dataloading, processing, and evaluation functionalities.
@@
    • maverix_<mode>.yaml: Individual setting files specifying the test settings for different modes. The currently supported modes are: audio_only, video_only, sub_only, video_sub, and video_audio.
    • Remark: Except for the open-sourced models listed in the paper or available through LMMs-Eval(Zhang et al. 2024; Li* et al. 2024), all other experiments require access to proprietary models.
@@
Appendix B B. Rationale and limitations
@@
Appendix C C. Additional experiments
@@
C.1 Standardized Prompt Design
To ensure fairness in evaluation, we adopt a standardized prompt format across all models. This eliminates potential biases introduced by prompt engineering and ensures that differences in performance are attributed to model capabilities rather than variations in input phrasing. For tasks requiring multiple modalities, the prompts incorporate the modalities in the inputs. In multiple-choice settings, the prompt is presented as follows: [images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available. [Subtitles] Select the best answer to the following multiple-choice question based on the video and the subtitles. Respond with only the letter (A, B, C, D, E, F, G, or H) of the correct option. [Question] The best answer is:
@@ C.1 Standardized Prompt Design
C.2 GPT-4o as Judge Criteria
For evaluating the open-ended answer qualities, we divide the evaluations in 4 different aspects: correctness, level of details, context understanding, and temporal consistency. We prompt the judge model as follows for correctness:
[…]
The template prompt for level of detail orientation:
[…]
The template prompt for level of context understanding:
[…]
C.3 Human Performance by Categories
We report performance across distinct agentic task categories in Tab. 7 to showcase the strengths and weaknesses of the human study participants within the defined contexts. We note that these statistics may be biased due to the composition of the participant pool, and further details on the recruitment process can be found in the Questionnaire Survey section.
@@ C.3 Human Performance by Categories
Visual Social 66.71 94.72 79.31
Sports 28.91 93.84 74.34
Information Querying 77.84 95.67 89.83
Sentiment 28.91 92.43 68.96
Shopping 51.62 94.11 83.37
Gaming 54.20 96.34 84.18
Overall 28.91 96.34 80
Audio Social 22.10 53.04 32.05
Sports 26.52 56.83 44.51
Information Querying 30.60 48.63 44.12
Sentiment 28.42 53.04 46.89
Shopping 30.60 79.56 58.48
Gaming 33.15 33.15 33.15
Overall 22.10 79.56 43.2
Both (A+V) Social 78 95.81 92.86
Sports 50.15 97.20 94.71
Information Querying 85.80 94.91 94.77
Sentiment 96.28 96.70 96.87
Shopping 70.20 95.17 93.60
Gaming 66.85 93.42 83.22
Overall 50.15 97.20 92.67
Table 7: Worst, Best, and Average human performance (%) across tasks and modalities.
C.4 Quantitative GPT Evaluations
To further assess model performance beyond multiple-choice accuracy, we conduct a quantitative evaluation of the open-ended question responses using GPT-based scoring. Specifically, as discussed in the paper, we prompt GPT-4o to act as an automated grader, evaluating the correctness and relevance of the model’s free-form answers based on predefined guidelines consistent with the task objectives in Tab. 8 and Tab. 9.
@@ C.4 Quantitative GPT Evaluations
We report the GPT-assigned scores across the full benchmark as well as per agentic task category, enabling a more granular understanding of the model’s strengths and weaknesses in open-ended scenarios. These findings complement the multiple-choice evaluations and expose areas where the model demonstrates strong language generation and areas where further improvement is needed. However, we note that these scores may be biased due to limitations in GPT-4o’s own reasoning and evaluation capabilities (Li et al. 2025a), which can introduce alignment artifacts or systematic preferences.
Model Temporal Consistency Correctness
Video Audio/Sub V + A/Sub Video Audio/Sub V + A/Sub
Human 1.85 1.69 2.83 1.93 1.46 2.65
Open-source MLLMs
EgoGPT-7B 1.3955 1.2523 1.5634 1.2477 1.1843 1.4930
Ola-7B 0.9178 1.2500 1.3638 0.9038 1.2171 1.3005
VITA 1.5 0.6643 1.2160 0.7042 0.6561 1.0587 0.7007
Qwen-2-VL 1.4554 1.3638 1.6772 1.3498 1.2254 1.6092
Qwen-2.5-VL 1.6819 1.2793 1.9730 1.4014 1.1279 1.7477
Qwen-2.5-Omni 1.6819 1.2793 1.9730 1.4014 1.1279 1.7477
InternVL2 0.9742 1.3169 1.1925 0.9237 0.9143 1.1315
LLaVA-OneVision 1.5962 1.3685 1.8052 1.4002 1.2899 1.5763
DeepSeekVL2-small 0.9507 1.3157 1.3286 0.9953 1.1866 1.3650
Proprietary MLLMs
Gemini 2.0-FL 1.6526 1.5880 2.1784 1.4296 1.4390 1.9014
Gemini 2.5-FL 1.7864 1.6244 2.1678 1.5751 1.4507 1.9167
Claude Sonnet 3.5 2.0798 1.9155 2.6009 1.6538 1.5857 2.2300
GPT-4o 1.5434 1.8462 2.4390 1.4167 1.5634 2.2207
GPT-4o-mini 1.9859 1.9272 2.3850 1.5739 1.6185 2.0117
Grok 4 2.2782 2.1854 2.8028 1.9507 1.7066 2.4354
NOVA-Lite 0.7453 0.9167 1.0129 0.9812 1.1068 1.2066
NOVA-Pro 0.9484 0.9930 1.3333 1.1901 1.1221 1.5387
Table 8: GPT-4o evaluation of Temporal Consistency and Correctness scores (higher is better). Columns show performance for Video, Audio/Subtitles, and their combination.
Model Context Detail Orientation
Video Audio/Sub V +A/Sub Video Audio/Sub V +A/Sub
Human 2.13 1.93 3.11 1.84 2.02 2.96
Open-source MLLMs
EgoGPT-7B 1.6420 1.5599 1.8732 1.2042 1.1244 1.3509
Ola-7B 1.3146 1.6491 1.7254 1.0047 1.2805 1.3345
VITA 1.5 0.9906 1.4918 1.0610 0.7465 1.2735 0.7958
Qwen-2-VL 1.7570 1.6549 2.0505 1.4343 1.4178 1.7430
Qwen-2.5-VL 1.9824 1.6197 2.2829 1.7336 1.3885 2.0669
InternVL2 1.3462 1.3521 1.5164 1.1303 1.0505 1.2523
LLaVA-OneVision 1.8662 1.7136 2.0129 1.5129 1.4284 1.7031
DeepSeekVL2-small 1.3920 1.6984 1.7418 1.0340 1.5998 1.3580
Proprietary MLLMs
Gemini 2.0-FL 1.8204 1.8439 2.1784 1.5880 1.5293 2.0117
Gemini 2.5-FL 1.9930 1.9120 2.3216 1.7653 1.6678 2.0692
Claude Sonnet 3.5 2.1725 2.0493 2.6960 2.0129 1.9096 2.5575
GPT-4o 1.7570 2.0634 2.6103 1.4178 1.7946 2.3169
GPT-4o-mini 2.1444 2.1397 2.5716 1.7277 1.8662 2.2336
Grok 4 2.5141 2.2488 2.9636 2.4660 2.1185 2.9566
NOVA-Lite 1.4178 1.5516 1.6843 1.1749 1.3451 1.3955
NOVA-Pro 1.7312 1.5387 2.0610 1.5282 1.3439 1.8204
Table 9: GPT-4o evaluation of Context score and Detail Orientation score (higher is better). Columns show performance for Video only, Audio/Subtitles only, and Video+Audio/Subtitles.
C.5 Quantitative Evaluations by Categories
To provide a comprehensive understanding of model performance across the diverse challenge dimensions of the benchmark, we report quantitative results for all probed models across each defined split. This includes evaluations broken down by agentic task categories, difficulty levels, audio characteristics, topics, and input modalities, as shown in Tab. 10.
@@ C.5 Quantitative Evaluations by Categories
GPT-4o Gemini 2.5 FL Qwen 2.5 Omni Nova Lite Nova Pro Ola-7B
Difficulty
Easy 66.7 59.8 52.1 55.0 62.5 57.4
Medium 67.4 54.9 48.8 52.4 54.9 54.1
Hard 49.7 43.0 45.2 38.4 43.0 41.1
Agentic Tasks
Social 62.6 55.6 47.5 48.5 50.5 52.5
Sentiment 56.8 54.3 58.0 42.0 55.6 55.6
Egocentric Agent 59.3 39.0 42.4 44.1 54.2 57.6
Information Querying 68.3 60.0 45.3 52.1 54.7 49.1
Sports 60.5 47.1 49.6 50.4 58.0 52.9
Gaming 61.2 49.5 41.7 39.8 47.6 46.6
Shopping 68.3 61.9 64.3 69.0 67.5 63.5
Audio Type
Natural Sound 52.9 51.5 52.9 35.3 45.6 48.5
Speech 66.4 56.7 47.8 53.8 57.6 52.0
Music 61.5 50.0 53.8 42.3 57.7 38.5
Artificial Sound 64.0 52.0 48.0 48.0 50.0 58.0
Mixed Sounds 63.0 53.1 51.5 51.5 56.1 56.5
Topics
Humanities and Society 71.4 64.7 56.4 54.1 63.9 54.1
Geography and Travel 56.2 50.0 37.5 56.2 31.2 56.2
Technology and Gaming 65.6 55.0 42.0 43.5 48.9 48.9
Science and Knowledge 60.7 50.0 50.0 46.4 35.7 60.7
Movies, TV and Animations 53.7 37.3 34.3 43.3 52.2 38.8
Arts and Performance 51.7 44.8 51.7 44.8 44.8 48.3
Pets and Animals 57.8 66.7 55.6 40.0 53.3 53.3
Business and Commerce 69.8 66.3 68.6 73.3 69.8 66.3
Life and Practical Skills 72.1 60.6 51.9 60.6 61.5 61.5
Sports and Adventure 61.1 49.2 48.4 46.0 58.7 50.0
Social Trends and Reactions 53.7 48.1 44.4 46.3 40.7 50.0
Vehicles and Transportation 66.7 33.3 55.6 44.4 77.8 66.7
Low Quality and Extended Content 58.3 29.2 25.0 41.7 50.0 37.5
Multimodal Abilities
Visual Reasoning 58.3 54.4 45.6 50.0 51.8 52.6
Emotional Inference 53.5 65.1 51.2 39.5 44.2 44.2
Situational Reasoning 70.0 65.0 35.0 55.0 65.0 70.0
Causal Reasoning 67.2 49.3 44.8 49.3 55.2 47.8
Factual Recall 67.3 55.4 52.9 53.8 58.6 54.2
Spatial Reasoning 55.2 41.4 48.3 34.5 55.2 37.9
Relationship 68.8 50.0 50.0 56.2 62.5 75.0
Temporal Distance 75.0 50.0 50.0 41.7 50.0 58.3
Overall 64.0 54.7 49.5 50.9 55.8 53.1
@@
Table 10: Performance comparison across different models and categories with Video + Audio/Subtitle modalities.
C.6 Common Error Modes